2. Use the following regression equation to answer the question:

Y= 1+3.98( X1) - 2.89 ( X2) -4.5( X3)

X1=distress score

X2=hours exercising

X3=counseling sessions attended

Y=minutes of ritualistic behavior

If a patient has a distress score of 5, spends 2 hours exercising and attends a counseling session you would expect to see about how many minutes of ritualistic behavior (Y)?

a. 29.5

b. 68.6

c. 10.6

d. Unable to determine
Answer #1

2)

Y= 1+3.98( X1) - 2.89 ( X2) -4.5( X3)

X1=distress score

X2=hours exercising

X3=counseling sessions attended

Now for x1=5

X2=2

X3= 1

Now y= 1+3.98(5)-2.89(2)-4.5(1)

= 1+19.9-5.78-4.5

= 10.62

Option c is correct.

Please like ?